OpenStudy (anonymous):

i need really slow instructions on how to find the distance formula between a pair of points the points are k(6,6) d(-3,-3)

OpenStudy (radar):

Are you familiar with the distance formula?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but i need some instructions on how to use it

OpenStudy (radar):

\[\sqrt{(x1-x2)^{2}+(y1-y2)^{2}}\]Is this the distance formula that you use?

OpenStudy (radar):

Substitute the x and y values from your two points k and d. Select one of the points as P1(x,y) and the other P(x,y). The selection is arbitrary, but once you select stay with it. In this case let point k be P1 and point d be P2. Any questions?

OpenStudy (radar):

P1 (6,6) x1=6 and y1=6. P2 (-3,-3) x2=-3 and y2=-3

OpenStudy (radar):

Hey is this slow enough?

OpenStudy (radar):

\[\sqrt{(6-(-3))^{2}+(6-(-3))^{2}}=\sqrt{9^{2}+9^{2}}=\sqrt{162}\]

OpenStudy (radar):

\[\sqrt{162}=9\sqrt{2}\]

OpenStudy (radar):

You agree Jim?

jimthompson5910 (jim_thompson5910):

sry was afk, yes that's correct

OpenStudy (radar):

Don't know what happened to the person asking the question?

jimthompson5910 (jim_thompson5910):

I'm sure they'll come back

OpenStudy (radar):

I wanted to attempt to have them plot the two points, and relate the distance formula to the triangle formed by dropping a vertical and a horizontal line showing that the distance formula and pythagorem theorem is the same thing.

jimthompson5910 (jim_thompson5910):

ah that's a good way to show how the distance formula works